OpenStudy (anonymous):

Please help! Permutations. Can someone help explain them? Please help! I have no idea what this is! It's very confusing! Thank you. :) For the questions simplify the expression. 1. 9! a. 45 b. 362,880 c. 265,456 d. 125,963 2. 25P3 a. 12,755 b. 14,290 c. 13,800 d. 18,250

OpenStudy (muhammad_nauman_umair):

k

OpenStudy (muhammad_nauman_umair):

this is example of how to open a factorial 4! = 4 × 3 × 2 × 1 = 24 now try to open the first one

OpenStudy (anonymous):

9!=\[\prod_{i=1}^{9} i\]

OpenStudy (anonymous):

Okay, so 9! = 9 x 8 x 7 x 6 x 5 x 4 x 3 x 2 x 1 = 362,880 Am I right?

OpenStudy (anonymous):

yes

OpenStudy (muhammad_nauman_umair):

yes

OpenStudy (anonymous):

Okay, thanks! Now how would I solve the next one? :)

OpenStudy (anonymous):

\[25P3=25C3*3!\]

OpenStudy (anonymous):

\[25C3=25C22\]

OpenStudy (anonymous):

so it is \[25C22*3!\]

OpenStudy (anonymous):

\[25*24*23=13800\]

OpenStudy (anonymous):

\[nCr=\frac{ n! }{ (n-r)!r! }\]

OpenStudy (anonymous):

Oh, that makes so much sense now! Thank you! :)

OpenStudy (muhammad_nauman_umair):

Simply u can also find permutation by this formula nPr=n!/(n−r)!